OpenStudy (anonymous):

MEDAL AND FAN...ASAP How many solutions are there for the pair of equations for lines Q and S? Explain your answer.

OpenStudy (anonymous):

A coordinate plane is shown with two lines graphed. Line Q has a slope of one half and crosses the y axis at 3. Line S has a slope of one half and crosses the y axis at negative 2.

OpenStudy (raymonde20):

What do you think it is?

OpenStudy (anonymous):

@Luv_luvi7 @Hayleymeyer

OpenStudy (anonymous):

Is there a picture?

OpenStudy (anonymous):

you dont need one

OpenStudy (anonymous):

no the pic would not come up but my first comment is the graph explained, that how it pasted @OM14forever

OpenStudy (anonymous):

because both lines have the same slope, they are parallel which means that have no solution at all because they don't intersect.

OpenStudy (welshfella):

line Q y = 0.5x + c when x = 0 y = 3 4 = 0.5*0 + c c = 3 equation of Q is y = 0.5x + 3

OpenStudy (anonymous):

there are no solutions at all because the two lines don't intersect. that is your answer.

OpenStudy (anonymous):

thank u @Luv_luvi7

OpenStudy (welshfella):

yes I was only the questioner how to find the equation of the line

OpenStudy (welshfella):

to find p (which is the fifth root of 243) use calculator enter 243 then '^' then (1/5) ( or 0.2)

OpenStudy (anonymous):

your welcome. there is no need to make it so complicated @welshfella